OpenStudy (anonymous):

Solve. 0.3x-0.5(2x+3)=-0.8(x+2)

OpenStudy (wolf1728):

Do the multiplications 0.3x-0.5(2x+3)=-0.8(x+2) .3x -x -1.5 = -.8x -1.6 Add up the terms .1x = -.1 Now all you have to do is solve for x